Define the term adiabatic lapse rate. obtain an expression for adiabatic lapse rate for the earth atmosphere.
1
Expert's answer
2020-08-18T11:16:32-0400

Lapse rate shows how much changes the temperature with height

"\Gamma = - \frac{dT}{dz}"

Even in condition of thermal equilibrium, there won't be equal temperature in different layer of gas due to Boltzmann distribution. If the the gas (e.g. in atmosphere) moves in a way that "PV^\gamma =" const, this processes is called adiabatic. In adiabatic atmosphere adiabatic lapse rate can be obtained.

From the first law of thermodynamics: "dU = \delta Q - PdV"

Since the processes are adiabatic, "\delta Q =0." Also, "dU = m c_V dT" .

From the equation of adiabatic process taking the differential: "V^\gamma dP + P \gamma V^{\gamma-1} dV =0 => PdV = -\frac{V^\gamma dP}{\gamma V^{\gamma-1}}"

"pdV = -\frac{VdP}{\gamma}"

Putting all result into the first law of thermodynamics, we get

"m c_V dT = \frac{VdP}{\gamma}"

Assuming an atmosphere in hydrostatic equilibrium "dP = -\rho g dz = -\frac{m g dz}{V}" and taking into account "\gamma = c_p/c_V" we have

"dT = -\frac{V/m * \rho g dz}{\gamma c_V}=-\frac{ g dz}{ c_P}"

"\Gamma = -\frac{dT}{dz}=\frac{g}{c_P}"